# Assessing the Reliability and Validity of GPT-4 in Annotating Emotion Appraisal Ratings

### v1

@@ 3.3 Prompt
Our prompt to the GPT-4 consists of three parts: 1) the general context C, 2) more detailed instructions I and 3) the description of the event D.
@@ 4.1 Q1 Reliability of GPT-4
Situat. Resp. 1.43 1.65 1.67 1.10 1.03 1.12 1.14 1.50
Own Resp. 1.34 1.45 1.60 1.03 0.83 1.06 0.91 2.26
Others’ Resp. 1.51 1.60 1.86 1.26 1.01 1.24 1.06 1.68
Anticip. Conseq. 1.60 1.69 1.68 1.32 1.21 1.35 1.28 1.57
@@
Situat. Control 1.58 1.65 1.66 1.10 0.99 1.12 1.14 1.10
Accept. Conseq. 1.65 1.84 1.64 1.25 1.39 1.27 1.43 1.39
@@
Macro avg. 1.45 1.57 1.61 1.12 0.99 1.15 1.06 1.22
Table 2: RMSE results for Q2, Q3, and Q4 (GPT-4 - GPT-4 annotator, Human - human reader-annotator, avg - average of five GPT-4 completions/human guesses, maj - majority vote of five GPT-4 completions/human guesses, conf - majority vote of five GPT-4 completions/human guesses with confidence as a tiebreaker, emo - majority vote of five GPT-4 completions with the emotion prediction task in a prompt)
@@ 4.3 Q3 Effect of Majority Voting
In previous work (Troiano et al., 2023), the accuracy of human reader-annotators guesses was considerably improved when their ratings by the annotators were aggregated using the majority voting over the guesses of five annotators. Encouraged by this, we used GPT-4 to generate five completions for each text in the reader-annotator dataset, and tested whether the majority voting has an impact on prediction accuracy. The ties were resolved based on the order of the ratings, which was essentially random, i.e., we chose the majority value, which was presented first in the dataset. In total, we generated predictions for 1200 $\times$ 5 data samples using a prompt I1 with added question "How confident is the model about chosen ratings?".
@@ 4.5 Q5 Impact of Event Description Length
5 Discussion and Conclusion
@@
We also tested the effect of predicting the emotion category on the accuracy of the appraisal ratings. Because different emotions are expected to have distinct appraisal profiles and thus might help in predicting the appraisal ratings, we did not observe improved results. This might be due to the accuracy of the emotion prediction task being moderate, at 55%, and incorrect emotion category prediction might have confused the model. We also observed that adding more complexity to a prompt resulted in less consistent responses and higher RMSE scores.
Ethical Considerations, Impact, and Limitations
Using GPT-4 in a reader-annotator context raises ethical questions regarding the accuracy and bias of the annotations. Therefore, GPT-4 annotations should always be reviewed and validated by human experts. In our research, we addressed this ethical consideration by only comparing GPT-4 with human annotations and not using event descriptions without relevant human reader-annotator ratings.
@@
In our research, GPT-4 predictions performed similarly to human reader-annotators in annotating appraisal ratings and thus could be applied in psychological research and practice. Predicting user appraisal profile of emotional events could help to identify behavioral and emotional patterns, support therapeutic interventions, or have other practical applications. However, in those settings, the requirement to submit sensitive and private user content to the GPT-4 API might not be desirable. An alternative would be to use GPT-4 predictions to augment the training data to improve the accuracy of smaller, local, models that could be subsequently applied in more sensitive data settings.
An important limitation of our research is the limited generalizability of results. We used only one model, GPT-4, and thus our results cannot be generalized to other available LLMs, although we anticipate that open-source models would probably show poorer results. Another major limitation is that GPT-4 and other LLM models can only act as reader-annotators and are limited in the experiencer-annotators roles. Thus, our results might have limited impact in psychological research aiming to study the role of appraisals in emotional experiences, as this subjective information can only be provided by human experiencers.

### v2

@@ 3.3 Prompt
Our prompt to the GPT-4 consists of three parts: the general context C, more detailed instructions I and the description of the event D.
@@ 4.1 Q1 Reliability of GPT-4
Situational Responsibility 1.43 1.65 1.67 1.10 1.03 1.12 1.14 1.50
Own Responsibility 1.34 1.45 1.60 1.03 0.83 1.06 0.91 2.26
Others’ Responsibility 1.51 1.60 1.86 1.26 1.01 1.24 1.06 1.68
Anticipated Consequences 1.60 1.69 1.68 1.32 1.21 1.35 1.28 1.57
@@
Situational Control 1.58 1.65 1.66 1.10 0.99 1.12 1.14 1.10
Accepted Consequences 1.65 1.84 1.64 1.25 1.39 1.27 1.43 1.39
@@
Macro average 1.45 1.57 1.61 1.12 0.99 1.15 1.06 1.22
Table 2: RMSE results for research questions Q2, Q3, and Q4. GPT-4 stands for the GPT-4 annotator, Human for the human reader-annotator, avg: average of five GPT-4 completions/human guesses, maj: majority vote of five GPT-4 completions/human guesses, conf: majority vote of five GPT-4 completions/human guesses with confidence rating as a tiebreaker, emo: majority vote of five GPT-4 completions with the emotion prediction task in a prompt. Bold/underline marks the appraisal dimensions that are consistently predicted better/worse than the macro average.
@@ 4.3 Q3 Effect of Majority Voting
In previous work (Troiano et al., 2023), the accuracy of human reader-annotators guesses was considerably improved when their ratings were aggregated using the majority voting over the guesses of five annotators. Encouraged by this, we used GPT-4 to generate five completions for each text in the reader-annotator dataset, and tested whether the majority voting has an impact on prediction accuracy. If there was a clear majority vote, that rating was selected as the final prediction. In case of ties (e.g., either two ratings were predicted twice or all five ratings were different), we adopted the following essentially random procedure. While each text had five independent predictions from the GPT-4 model, we pretended to have done just a single completion and chose the rating from the first of the five runs. In total, we generated predictions for 1200 $\times$ 5 data samples using a prompt I1 with added question "How confident is the model about chosen ratings?".
@@ 4.5 Q5 Impact of Event Description Length
5 Discussion
[…]
In contrast to our expectations, asking the GPT-4 model to predict the emotion first and then the appraisal ratings (Q4) did not improve the overall prediction, although the emotion label was expected give useful information about the appraisal values. This might be due to the accuracy of the emotion prediction task being moderate, only at 55%, and incorrect emotion category predictions may have confused the model. To check that, we split the data into two groups by correctly and incorrectly predicted emotion labels. Although texts with correctly predicted labels demonstrated better macro average RMSE compared to the texts with incorrect labels (1.21 vs 1.24 RMSE), the difference is not large. However, a closer look at individual appraisal dimensions revealed that most appraisals showed significantly better or comparable accuracy for correctly predicted emotions, with few exceptions of Situational Responsibility (1.57 vs 1.42), Others’ Responsibility (1.70 vs 1.65), Anticipated Consequences (1.65 vs 1.48), and Not Consider (1.25 vs 1.10). A more fine-grained analysis of why the predictions of these dimensions were more accurate with incorrect emotion labels remains for future research.
We also found that shorter event descriptions generally exhibit lower RMSE values. This result is expected as very short texts convey too little information to predict various appraisal aspects accurately. We found that the optimal event description length that appears to start from the range of roughly 400 to 500 characters and although the improvement in prediction accuracy is more steep for the GPT-4 model, a similar pattern can be observed also for human-readers. This finding has implications for collecting data for emotion appraisal research, as it suggests that researchers should aim for eliciting event descriptions at least 500 characters long. Finally, throughout the research, we observed that adding more complexity to a prompt resulted in less consistent responses as well as higher RMSE scores. This observation is in line with the reports of other researchers (Herderich et al., 2024).
6 Conclusions
@@
Impact
In our research, GPT-4 predictions performed similarly to human reader-annotators in annotating appraisal ratings and thus could be applied in psychological research and practice. Predicting user appraisal profile of emotional events could help to identify behavioral and emotional patterns, support therapeutic interventions, or have other practical applications.
Our study also contributes to the set of effective strategies in predicting appraisal ratings by GPT-4 or potentially similar LLMs. We show that adopting majority voting algorithm based on five completions can considerably improve the performance of this subjective task. Moreover, we empirically establish an minimum optimal event description length below which both human readers and GPT-4 model prediction accuracy starts to degrade, thus providing a practical guideline for appraisal researchers interested in using predictive models.
The results of this work can inform further research in developing automated reappraisal self-help systems or similar applications, offering practical tools for emotional reframing and appraisal annotation. However, in those settings, the requirement to submit sensitive and private user content to the GPT-4 API might not be desirable. An alternative would be to use GPT-4 predictions to augment the training data to improve the accuracy of smaller, local, models that could be subsequently applied in more sensitive data settings.
Limitations
An important limitation of our research is the limited generalizability of the results. We used only one model, GPT-4, and thus our results cannot be generalized to other available LLMs, although we anticipate that open-source models would probably show poorer results. We also used the specific default configuration of the GPT-4 model. LLMs are sensitive to hyperparameters tuning, and our findings can be applicable only to the used settings.
Another significant limitation arises from the dataset characteristics. The dataset used in this study reflects limited demographic and socioeconomic diversity, and the emotional events reported by crowd-source participants may be synthetic, potentially reducing their relevance to real-world contexts.
Finally, GPT-4 and other LLM models can only act as reader-annotators and are limited in the experiencer-annotators roles. Thus, our results might have limited impact in psychological research aiming to study the role of appraisals in emotional experiences, as this subjective information can only be provided by human experiencers.
Ethical Considerations
Using GPT-4 in a reader-annotator context raises ethical questions regarding the accuracy and bias of the annotations. Therefore, GPT-4 annotations should always be validated by both actual event experiencers and human reader-annotators.
@@
However, the use this dataset raises some considerations regarding its intended scope and limitations. The dataset was crowd-sourced and as such, we can assume that the participants were a sample from a generally healthy population. This constraint needs to be kept in mind when using the results of our research in designing systems for clinical domains, such as self-help systems intended to aid in emotional reappraisal for people with clinical issues.